An examination paper consists of 5 questions in section A and 5 questions in section B. A total of 8 questions must be answered. In how many ways can a student select the questions if he is to answer at least 4 questions from section A.


1
Expert's answer
2021-07-06T03:40:48-0400

Student has to solve at least "4" questions from section A.

So, he can choose in 2 ways :

"a)" "4" from section A and "4" from section B.

"b)" "5" from section A and "3" from section B.

Total number of ways of selecting the questions "N=" "N_a+N_b"

"\ N=^5C_4\times ^5C_4 + ^5C_5\times ^5C_3"

"N=\frac{5!}{(5-4)!(4!)}\times \frac{5!}{(5-4)!(4!)}+\frac{5!}{(5-5)!(5!)}\times \frac{5!}{(5-3)!(3!)}"

"N=\frac{5!}{1!4!}\frac{5!}{1!4!}+\frac{5!}{0!5!}\frac{5!}{2!3!}=5\times 5+ 1\times 10"

"N=25+10=35"